OpenStudy (mathmath333):

the difference between two numbers is 16.what can be said about the total numbers divisible by 7 that lie between these two numbers.

ganeshie8 (ganeshie8):

what do you think ? for example, consider the numbers between 0 and 16

OpenStudy (mathmath333):

0,7,14.. total 3 no's

OpenStudy (aum):

Let 'a' and 'b' be the two numbers, with 'a' being the smaller number and 'b' being the bigger number and b = a + 16 Consider three cases: 1) 'a' is divisible by 7. Example: a =14 and b = 30 2) 'b' is divisible by 7. Example: a = 5 and b = 21 3) Neither 'a' nor 'b' is divisible by 7. Example: a =13 and b = 29 Find the total number of numbers divisible by 7 in each interval that includes both end values.

OpenStudy (aum):

Note: We did not consider the fourth case of BOTH 'a' and 'b' being divisible by 7 because that is not possible when the difference between the two numbers, which is 16, is not divisible by 7.

OpenStudy (mathmath333):

thnks

OpenStudy (anonymous):

only one number or many numbers which lie in the interval

OpenStudy (aum):

Well, if you consider the case of a = 15 and b = 31 we have only 21 and 28 divisible by 7.

OpenStudy (mathmath333):

in first two cases the total numbers divisible by 7 are 3.. but in the third case it varies from 2,3.

OpenStudy (aum):

So minimum 2 and maximum 3.

OpenStudy (mathmath333):

yes

OpenStudy (abb0t):

Nicely explained @aum unlike some other users who ask rhetorical questions from the get-go.

OpenStudy (aum):

Thank you abb0t.

OpenStudy (mathmath333):

am i correct ?

ganeshie8 (ganeshie8):

see if below helps to conclude : There is exactly one and only one number divisible ny \(n\) in \(n\) consecutive integers.

ganeshie8 (ganeshie8):

@BSwan wanna try the proof for above statement ? :)

OpenStudy (anonymous):

i wanna lol brb , im washing clothes

ganeshie8 (ganeshie8):

most proofs are done while thinking in shower or wash room :P I thihnk this proof involves an application of DA algorithm - it could be more involved too... not sure

ganeshie8 (ganeshie8):

you need to prove both existence and uniqueness

OpenStudy (anonymous):

could you rewrite the statment plz ?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

rewrite ?

OpenStudy (anonymous):

yeah :D

ganeshie8 (ganeshie8):

take any "n" consecutive integers, prove that one number will always be divisible by "n"

OpenStudy (anonymous):

ok , i think we proved that before wilson thm , right ?

ganeshie8 (ganeshie8):

for example : take 5 consecutive numbers {7,8,9,10, 11} 5 | 10

ganeshie8 (ganeshie8):

i don't remember wilson's proof need to look up again

OpenStudy (anonymous):

so let a be the first number that devide n , then a+1=1 mod n a+2=2 mod n a+3=3 mod n . . . . a+(n-1)= n-1 mod n a+n=0 mod n and ect :)

OpenStudy (anonymous):

@ganeshie8 any other method ?

OpenStudy (anonymous):

\(\Large\bf\rlap{\color{red}{\bigstar~T}\color{gold}{H}\color{greenyellow}{A}\color{lime}{N}\color{cyan}{K}\color{blue}{ ~}\hspace{0 pt}\color{purple}{Y}\color{magenta}{O}\color{orchid}{U~}\color{red}{S}\color{gold}{O~}\color{lime}{M}\color{cyan}{U}\color{blue}{C}\color{blueviolet}{H}\color{purple}{ }\color{magenta}{ }\color{orchid}{ }\color{pink}{ }\color{red}{!~\bigstar}}{\color{black}{\; \bigstar~THANK ~YOU ~SO~MUCH!~\bigstar}} \)